DIODE LASER SPECTRUM OF JET-COOLED CH_{2}F_{2} IN THE 9-10 μm REGION

\underline{JANINA MATUSZESKI} and MARK D. MARSHALL, Department of Chemistry-2243, Amherst College, P.O. Box 5000, Amherst, MA 01002-5000.

Utilizing CO_{2} laser optical pumping in the 9-10 μm region, methylene fluoride is a common gain medium in far-infrared gas lasers. Unfortunately, this part of methylene fluoride’s infrared spectrum is very dense and not well characterized, making it difficult to identify the transitions responsible for FIR laser action. Four vibrational bands (ν_{3}, ν_{7}, ν_{9}, 2ν_{4}) are present with several Coriolis coupled to one another. Hot bands and high-J transitions contribute to the congestion. Rotational and vibrational cooling of methylene fluoride in a pulsed-jet expansion leads to a much simpler spectrum, which we have obtained using a tunable diode laser spectrometer. Work on the ν_{3}, ν_{7}, and ν_{9} bands has been completed and Coriolis effects have been studied.